# A Fine-tuning Dataset and Benchmark for Large Language Models for Protein Understanding

### v1

@@ Abstract
The high similarities between protein sequences and natural language, particularly in their sequential data structures, have driven parallel advancements in deep learning models for both domains. In natural language processing (NLP), large language models (LLMs) have achieved remarkable success in tasks such as text generation, translation, and conversational agents, owing to their extensive training on diverse datasets that enable them to capture complex language patterns and generate human-like text. Inspired by these advancements, researchers have attempted to adapt LLMs for protein understanding by integrating a protein sequence encoder with a pre-trained LLM, following designs like LLaVa. However, this adaptation raises a fundamental question: “Can LLMs, originally designed for NLP, effectively comprehend protein sequences as a form of language?” Current datasets fall short in addressing this question due to the lack of a direct correlation between protein sequences and corresponding text descriptions, limiting the ability to train and evaluate LLMs for protein understanding effectively. To bridge this gap, we introduce ProteinLMDataset, a dataset specifically designed for further self-supervised pretraining and supervised fine-tuning (SFT) of LLMs to enhance their capability for protein sequence comprehension. Specifically, ProteinLMDataset includes 17.46 billion tokens for pretraining and 893K instructions for SFT. Additionally, we present ProteinLMBench, the first benchmark dataset consisting of 944 manually verified multiple-choice questions for assessing the protein understanding capabilities of LLMs. ProteinLMBench incorporates protein-related details and sequences in multiple languages, establishing a new standard for evaluating LLMs’ abilities in protein comprehension. The large language model InternLM2-7B, pretrained and fine-tuned on the ProteinLMDataset, outperforms GPT-4 on ProteinLMBench, achieving the highest accuracy score. The dataset and the benchmark are available at https://huggingface.co/datasets/tsynbio/ProteinLMBench.
@@ 4.4 Text input filtering and tokenization procedure
Tokens statistic of the fine-tuning dataset
Segment Scope Min Len. Average Len. Max Len.
Enzyme CoT molecule’s expression 84 800 15300
UniProt Function protein functionality 74 3610 70500
UniProt Induction protein induction 82 1650 31300
UniProt Involvement in disease protein diseases 88 3495 68800
UniProt Sub-unit structure protein structure 74 3560 70500
UniProt Tissue specificity tissue gene expression 65 145 2310
@@ 5 Quality of dataset and benchmark
(a) Statistics for all tokens
(b) Statistics for tokens from UniProtKB and Seq/PubMed
@@
Diversity and statistics of the ProteinLMBench
@@
Safety during collection
@@ C.1 LLMs instruction
We conducted three different experiments: no training, supervised fine-tuning (SFT), and self-supervised learning combined with supervised fine-tuning (SSL-SFT) using the ProteinLMDataset. For the no-training experiment, we employed various large language models, including Falcon-7B [2], Qwen1.5-7B [3], Moonshot [36], Mistral-7B-Instruct-v0.2 [14], Baichuan2-7B-Chat [4], Llama-2-7B-Chat-hf [30], InternLM-Chat-20B, InternLM2-Chat-7B, and its variations [7], ChatGLM3-6B [9], Yi-6B-Chat [1], GPT3.5-turbo, and GPT4.0-turbo [20]. For the SFT and SSL-SFT experiments, we utilized the InternLM2-7B model, naming them InternLM2-Protein-7B (w/o SSL) and InternLM2-Protein-7B, respectively. The accuracy of all models was evaluated using our proposed ProteinLMBench benchmark.
C.1.1 Hyperparameter configuration
@@
C.1.2 Self-Supervised learning (SSL) phase
For the self-supervised learning (SSL) phase, we initialized the model weights based on the InternLM2 model. We trained on a mixed dataset comprising 50 billion tokens, integrating both our specific protein data and a broad range of general-purpose data. The training was performed with a batch size of 16 and a fixed learning rate of 0.0002. The same SSL process, involving initialization from a pre-trained model and training on a mixed dataset, was followed for the other large language models (LLMs) as highlighted in the main manuscript.
C.1.3 Supervised fine-tuning (SFT) phase
@@
Table 3: Accuracy and inference time of protein language models on ProteinLMBench.
@@
C.2 Model performance comparison
Tab. 3 compares the accuracy and inference time of various protein language models on the ProteinLMBench benchmark. GPT4.0-turbo achieved the highest accuracy of 57.94% with a relatively fast inference time of 15.52 minutes among models without fine-tuning on ProteinLMDataset. InternLM2-20B and GPT3.5-turbo also showed competitive accuracies of 57.52% and 55.19%, respectively. Notably, InternLM2-Protein-7B, which utilized self-supervised learning on the ProteinLMDataset followed by supervised fine-tuning, achieved the highest accuracy of 62.18% among all models, with a reasonable inference time of 22.34 minutes. Models like Falcon-7B, Qwen1.5-7B, and Moonshot exhibited lower accuracies, while others like Yi-6B, InternLM-Chat-20B, and Llama2-7B had considerably longer inference times. The results highlight the trade-off between accuracy and inference time, with InternLM2-Protein-7B demonstrating the best balance of high accuracy and reasonable inference time.
Appendix D Thoroughly related work
Existing Chinese-English datasets. Traditional benchmarks in the English language primarily focus on evaluating specific Language Model (LLM) abilities related to singular tasks, such as providing answers in general question answering, text summarization, or general linguistic understanding [31, 22, 18].
Recent benchmarks, such as MMLU [12], expand the evaluation scope by encompassing multi-domain and multi-task assessments derived from real-world university and high school exams and published manuscripts. BIG-bench [27] comprises 204 diverse tasks, including some considered beyond current LLM capabilities, while HELM [17] aggregates 42 tasks evaluating LLM performance across seven different metrics, ranging from accuracy to robustness. In the domain of NLP benchmarks specific to the Chinese language, two notable studies have made significant strides. The first study, as outlined in [13], is remarkable for its thorough evaluation of various Language Model (LLM) systems. Notably, GPT-4 achieves an impressive accuracy exceeding 60%. The dataset comprises 13,948 questions and answers, spanning diverse proficiency levels (middle school, high school, college, professional) and covering a wide array of topics including STEM, Humanities, Social Science, and more. Despite its merits, the evaluation has certain limitations. It is confined to regions within China, predominantly focuses on accuracy metrics, lacks detailed distribution statistics, and falls short in offering interpretability and insights conducive to enhanced reasoning. While the inclusion of zero-shot and five-shot comparisons enriches the assessment, there remains room for improvement in terms of comprehensiveness and interpretive depth. Conversely, the second study, detailed in [35], caters specifically to the needs of native Chinese readers proficient in both modern and classical Chinese. The dataset encompasses 8,390 documents, each containing 1,024 characters, spanning classic literature, modern articles, and classical poetry. With 20,477 questions of varying language difficulty, it incorporates a rich variety of characters in each document style, facilitating diverse language exploration. The benchmark, featuring models that achieve an impressive 79% accuracy, stands out for its provision of detailed statistics on dataset distribution, ensuring a fair evaluation. However, a limitation of this study lies in the absence of scientific documents, representing a unique writing style not covered in this benchmark. This gap highlights the need for further exploration and inclusion of diverse textual genres to provide a more comprehensive assessment.
The existing literature on English benchmarks primarily focuses on specific model abilities for singular linguistic tasks, with limited extension to multi-domain and multi-task assessments. However, in the context of Chinese NLP benchmarks, significant limitations emerge, including rational biases with geographic restrictions to China regions dataset collection, absence of analytical documentation and statistics, and lack of interpretability with limited insights into reasoning abilities. Overall, these limitations underscore the need for more comprehensive and diverse multi-task language datasets, addressing a broader geographic scope of collection, incorporating varied evaluation metrics, and ensuring coverage of distinct writing styles, including scientific content. Notably, there is currently no dataset known to the authors that provides text-to-text pairs for English and Chinese translation.
Existing protein sequence datasets. The UniProtKB [8] provides a comprehensive repository of protein sequences along with functional annotations. With a vast collection exceeding 227 million sequences, it serves as a vital resource for a wide range of research applications. Ongoing efforts are directed towards establishing a reference proteome for each taxonomic group, ensuring inclusivity across diverse organisms. Continuous literature extraction mechanisms ensure the entries remain current, while unreviewed entries benefit from automated system annotations facilitated by machine learning techniques, thereby enhancing overall coverage. UniProt demonstrates remarkable adaptability to the evolving information landscape by continually enhancing its data processing pipeline and website interface. However, it is imperative to acknowledge potential biases inherent in community contributions, which may affect the representativeness of the dataset. Moreover, the utilization of automated systems introduces the possibility of errors, thereby impacting the accuracy of annotations.
Annotating proteins that are yet to be characterized presents inherent challenges, and limitations persist in achieving comprehensive characterization. Efficient management of the influx of whole genome sequencing data poses significant challenges, and integrating diverse external resources may complicate the coherent presentation of data. The PEER dataset [33] fills a crucial gap by providing a standardized benchmark for evaluating deep learning methods in protein sequence understanding. Its openly available datasets and source codes enhance transparency and reproducibility in research. Covering diverse tasks like function prediction, localization, structure, protein-protein, and protein-ligand interactions, PEER evaluates various sequence-based methods, with a focus on the effectiveness of large-scale pre-trained protein language models, notably ESM-1b. Multi-task learning significantly boosts performance across tasks. However, extending beyond sequence-based datasets presents challenges, necessitating ongoing community collaboration to ensure sustained effectiveness.
Several important datasets related to protein sequences include RefSeq [19], PDB [6], CATH (Class, Architecture, Topology, Homology) [21], and CASP15 [26, 16]. RefSeq provides a comprehensive, curated collection of reference sequences for various organisms, though it primarily focuses on single-chain protein design. CATH, while valuable, has limitations in representing the full diversity of protein structures [6]. Incorporating de novo proteins from CASP15 [26, 16], the TS45 dataset comprises diverse protein targets, including Free Modeling (FM) and Template-Based Modeling (TBM), divided into TBM-easy and TBM-hard proteins. TS45 serves as a benchmark for de novo protein design, offering dissimilarity to known structures and mitigating biases present in training sets. Lastly, the study by [10] introduces a biomolecular instruction dataset aimed at enhancing LLMs’ understanding and prediction capabilities in this specialized field. While promising, it’s important to note that this benchmark is still in its early stages and has limited applicability for real-world tasks. Further research is required to optimize instruction tuning methods, refine the model’s ability to understand and follow instructions, and improve its generalization across various protein and biomolecular tasks.
Protein design databases. KEGG [15] represents a curated resource, offering researchers a comprehensive repository of accurately annotated biological entities, spanning molecular networks, genes, proteins, chemicals, and health-related data. Despite its strengths, KEGG faces limitations such as a current focus on disease-related perturbations in human cells, potential biases in data inclusion, and the resource-intensive nature of manual curation. Meanwhile, the STRING database [29] distinguishes itself through its extensive integration of known and predicted protein associations, encompassing both physical interactions and functional relationships across over 14,000 organisms. Despite its comprehensive coverage, STRING encounters challenges including potential biases in text mining, reliance on predictive methods, and the absence of options for tissue-specific network pruning. Lastly, InterPro [5], a leading protein classification database, offers comprehensive annotations through the integration of predictive models, providing insights into families, domains, and conserved sites. However, its limitations include a lack of updates, interpretability, and sustainability.
While the protein-related databases and datasets mentioned offer invaluable resources for various aspects of biological research, they are not without their limitations. These constraints include potential biases stemming from community contributions and automated systems in UniProtKB, challenges in extending beyond sequence-based datasets in PEER, a focus solely on single-chain structures in CATH, limitations in representing the complete diversity of protein structures in PDB, applicability primarily to designs present in training sets in CASP15, a disease-centric focus and resource-intensive manual curation in KEGG, potential biases in text mining and reliance on predictive methods in STRING, and challenges related to funding growth and dependencies on individual resources in InterPro. Recognizing these limitations is crucial for researchers to navigate and interpret the data effectively and highlights the ongoing efforts needed to improve the comprehensiveness and sustainability of these resources.

### v2

@@ Abstract
The high similarities between protein sequences and natural language, particularly in their sequential data structures, have driven parallel advancements in deep learning models for both domains. In natural language processing (NLP), large language models (LLMs) have achieved remarkable success in tasks such as text generation, translation, and conversational agents, owing to their extensive training on diverse datasets that enable them to capture complex language patterns and generate human-like text. Inspired by these advancements, researchers have attempted to adapt LLMs for protein understanding by integrating a protein sequence encoder with a pre-trained LLM, following designs like LLaVa. However, this adaptation raises a fundamental question: “Can LLMs, originally designed for NLP, effectively comprehend protein sequences as a form of language?” Current datasets fall short in addressing this question due to the lack of a direct correlation between protein sequences and corresponding text descriptions, limiting the ability to train and evaluate LLMs for protein understanding effectively. To bridge this gap, we introduce ProteinLMDataset, a dataset specifically designed for further self-supervised pretraining and supervised fine-tuning (SFT) of LLMs to enhance their capability for protein sequence comprehension. Specifically, ProteinLMDataset includes 17.46 billion tokens for pretraining and 893K instructions for SFT. Additionally, we present ProteinLMBench, the first benchmark dataset consisting of 944 manually verified multiple-choice questions for assessing the protein understanding capabilities of LLMs. ProteinLMBench incorporates protein-related details and sequences in multiple languages, establishing a new standard for evaluating LLMs’ abilities in protein comprehension. The large language model InternLM2-7B, pretrained and fine-tuned on the ProteinLMDataset, outperforms GPT-4 on ProteinLMBench, achieving the highest accuracy score. The dataset and the benchmark are available at https://huggingface.co/datasets/tsynbio/ProteinLMDataset/ and https://huggingface.co/datasets/tsynbio/ProteinLMBench. The code is available at https://github.com/tsynbio/ProteinLMDataset/.
@@ 5 Quality of Dataset and Benchmark
a. Refer to caption
b. Refer to caption
@@
Tokens statistic of the fine-tuning dataset
Segment Scope Min Len. Average Len. Max Len.
Enzyme CoT molecule’s expression 84 800 15300
UniProt Function protein functionality 74 3610 70500
UniProt Induction protein induction 82 1650 31300
UniProt Involvement in disease protein diseases 88 3495 68800
UniProt Sub-unit structure protein structure 74 3560 70500
UniProt Tissue specificity tissue gene expression 65 145 2310
Diversity and statistics of the ProteinLMBench.
@@
Safety during collection.
@@ 7 Conclusion
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] Yes, they do.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 5.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A] Our benchmark and dataset will not lead to any potential negative societal impacts
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] We confirm that our paper conforms to them.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A] Our benchmark and dataset paper does not include any theoretical results.
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A] Our benchmark and dataset paper does not include any theoretical results.
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes]
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Appendix Section C.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] We did not conduct multiple experiments or set random seeds.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix Section C.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] See Section 2.
    2. (b)
      Did you mention the license of the assets? [Yes] We provide the license in the dataset hugging face link.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] See Section 2.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A] Our dataset does not contain any personal information.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Appendix Section B.2
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A] We haven’t used crowdsourcing or conducted research with human subjects.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A] We haven’t used crowdsourcing or conducted research with human subjects.
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A] We haven’t used crowdsourcing or conducted research with human subjects.
@@ C.1 LLMs Instruction
TourSynbio is a model developed by building upon the InternLM2 through further pre-training and supervised fine-tuning. It retains the original architectural design without any modifications.
C.2 Self-supervised training & Supervised fine-tune training details
C.2.1 Hyperparameter Configuration
@@
C.2.2 Self-Supervised Learning (SSL) Phase
For the SSL phase, we initialized our model weights based on the InternLM2 model. We trained on a mixed dataset comprising 50 billion tokens, integrating both our specific data and a broad range of general-purpose data. The training utilized a batch size of 16 and a fixed learning rate of 0.0002. The same process followed for the other LLM models is highlighted in the main manuscript.
C.2.3 Supervised Fine-Tuning (SFT) Phase
@@
C.3 Model performance comparison & Further analysis
Tab. 3 presents the analytical accuracy scores of the different LLM models for the self-supervise training (SSL) and the supervised fine-tuning training (SFT) experiments.
Table 3: Different protein language models accuracy score and inference time in the evaluation cohort.